OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates (-3, 5pi/3)

OpenStudy (loser66):

you have r and \(\theta\), just apply x = r cos \(\theta\) and y = r sin\(\theta\) why is it hard?

OpenStudy (anonymous):

i tried that but it kept giving me (-2.9875, -0.2738) for an answer but the choices that i have are in fractions.

OpenStudy (loser66):

so, convert it into fraction, that's it. use calculator.

OpenStudy (loser66):

my calculator said that x = -3/2, you calculate y

OpenStudy (loser66):

and you can "cheat" by look at the choices, which has x = -3/2 , no matter what y is, pick it, hehehe..